OpenStudy (anonymous):

r^3 - 4r +3 =0, factor this equation

OpenStudy (anonymous):

\[r^{3}-4r +3=0\]

OpenStudy (anonymous):

\[(r-1) \left(r^2+r-3\right)=0 \]

OpenStudy (anonymous):

can you do a step by step?

Directrix (directrix):

Do you know synthetic division and the rational roots theorem?

OpenStudy (anonymous):

The following might help. http://www.youtube.com/watch?v=mXdAiJY7YfY I used the Mathematica program for this problem.

OpenStudy (anonymous):

forgot a zero here for a place holder for the r^2 term. that factorization is not right